# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

### v3

@@ Abstract
The rapid advancement of large language models (LLMs) has revolutionized role-playing, enabling the development of general role-playing models. However, current role-playing training has two significant issues: (I) Using a predefined role profile to prompt dialogue training for specific scenarios usually leads to inconsistencies and even conflicts between the dialogue and the profile, resulting in training biases. (II) The model learns to imitate the role based solely on the profile, neglecting profile-dialogue alignment at the sentence level. In this work, we propose a simple yet effective framework called Beyond Dialogue, designed to overcome these hurdles. This framework innovatively introduces “beyond dialogue” tasks to align dialogue with profile traits based on each specific scenario, thereby eliminating biases during training. Furthermore, by adopting an innovative prompting mechanism that generates reasoning outcomes for training, the framework allows the model to achieve fine-grained alignment between profile and dialogue at the sentence level. The aforementioned methods are fully automated and low-cost. Additionally, the integration of automated dialogue and objective evaluation methods forms a comprehensive framework, paving the way for general role-playing. Experimental results demonstrate that our model excels in adhering to and reflecting various dimensions of role profiles, outperforming most proprietary general and specialized role-playing baselines. All code and datasets are available at https://github.com/yuyouyu32/BeyondDialogue.
Introduction
The rapid advancement of large language models (LLMs) has demonstrated their significant potential in user interactions (Achiam et al. 2023; Bai, Bai, and et al. 2023; Dubey et al. 2024). A particularly promising area is the development of role-playing LLM agents, which are capable of simulating both real and fictional roles to deliver immersive and interactive experiences (Chen et al. 2024b, a). Applications like Character AI (Shazeer and Freitas 2024), Replika (Replika 2024), and Baichuan-NPC (Baichuan 2024) exemplify the growing impact of these agents. However, the proprietary nature of these platforms restricts users’ ability to customize specific role models. Thus, the development of an open-source, general role-playing model is imperative. Such a model provides a reliable foundation for expressing predefined role profiles and customizing role models across scenarios (Tencent 2024; Baichuan 2024).
@@ Introduction
At the current stage of role-playing dialogue training tasks, role profiles are typically manually compiled (Shao et al. 2023) or generated through LLM summaries (Zhou et al. 2023), including information such as personality and speaking style as depicted throughout a whole novel or script (Zhou et al. 2023; Li et al. 2023). However, a bias exists between the training dialogues extracted from individual scenarios and the predefined role profiles. As an illustration in Figure 1, Hermione’s speaking style is predefined into four categories, which are manually extracted from the entire content of the novel. However, the training dialogues for the scenario in Figure 1 predominantly reflect Hermione’s bookish (blue) and encouraging (green) styles. This bias arises because the predefined profile is generally not fully represented in a single scenario, leading to inconsistencies where the prompted role does not align with, or even contradicts, the training dialogues. Such biases are common: in the HPD dataset (Chen et al. 2023), we sampled 800 scenarios and found that 69.68% of the dialogues showed a bias from the predefined character profiles, while 50.69% exhibited a bias in style representation. These inconsistencies between predefined role profiles and the training dialogue corpus, as notably evident in our §Main Results experiments, can mislead the model, hindering its ability to accurately reflect the role profile during dialogues. Meanwhile, the model has only learned a vague mapping from profile to dialogue, without fine-grained alignment of specific traits to their corresponding dialogues. As shown in Figure 1, the model fails to map the traits “bookish” and “encouraging” to the specific dialogue sentences highlighted in blue and green. Consequently, the model is unable to learn how “personality traits are revealed in dialogues”.
@@
We propose a new training role-playing framework—Beyond Dialogue to overcome the aforementioned issues and advance toward general role-playing. We incorporate powerful LLM through a prompting mechanism approach to align role profiles with scenario role-playing dialogues. This not only ensures the reliability of the LLM’s reasoning but also explicitly creates a fine-grained alignment dataset at the sentence level. Taking inspiration from actors learning to play different roles—understanding the performance of various role traits in scenarios to enhance their portrayal (Zarrilli 2005)—we also employ fine-grained alignment tasks to train the role-playing model. These beyond-dialogue alignment tasks help the model understand how to represent specific role traits at the sentence level, thereby enhancing its overall effectiveness in role portrayal.
During the evaluation process, we assess the model’s role-playing capability by examining how well the dialogue aligns with the defined role profiles, which is consistent with the beyond-dialogue alignment tasks. Traditional evaluation methods have largely depended on subjective assessments, such as human evaluations or LLMs providing consistent scores between dialogue and profiles. These approaches are often irreproducible and lead to significant evaluation variance. Consequently, we have converted all evaluation tasks into objective tasks, with the evaluation objective clearly defined as the model’s ability to follow the role profiles. To ensure both efficiency and reproducibility in the evaluation, we employ an approach that combines automatic dialogues with the “LLMs as Judger” (Kim et al. 2023) method to assess different role-playing models.
We applied Beyond Dialogue to bilingual chat LLM baselines—Qwen2 (Bai, Bai, and et al. 2023) and Mistral (Jiang et al. 2023)—to empirically evaluate its effectiveness. Extensive experiments demonstrate that Beyond Dialogue significantly enhances the ability of LLMs to follow and portray role profiles. Under our framework, the trained models surpassed existing proprietary advanced general baselines, such as GPT-4o (OpenAI 2024), and specialized proprietary role-playing baselines, like Baichuan-NPC-Turbo (Baichuan 2024), in key dimensions of general role-playing performance. Moreover, the beyond dialogue alignment tasks consistently improved the models’ role-playing evaluation accuracy by over 30%, approaching the evaluation capabilities of GPT-4o. This alignment supports the development of a self-generating, self-evaluating, and self-aligning pipeline for role-playing models, fostering continuous improvement in their performance.
@@
    We identified a bias in role-playing training where the predefined role’s profile does not align with the dialogues at the specific scenario scope. Furthermore, we demonstrated that this bias significantly hinders the model’s ability to effectively follow the predefined profile.
@@
    We propose Beyond Dialogue, a novel general role-playing training framework that employs LLMs with a prompting mechanism to align role profiles with scenario-specific dialogues. This approach derives fine-grained alignment tasks at the sentence level, which are incorporated into the role-playing training, further enhancing the effectiveness of role portrayal.
@@
    We introduce a novel evaluation pipeline that converts all assessment tasks into objective tasks, focusing on the model’s ability to follow user-defined role profiles. By combining automatic dialogues with the “LLMs as Judger” method, our approach ensures both efficiency and reproducibility in role-playing model evaluation.
Related Work
@@
LIGHT (Urbanek et al. 2019) Sentence
PDP (Han et al. 2022) Sentence
ChatHaruhi (Li et al. 2023) Sentence
HPD (Chen et al. 2023) Sentence
CharacterLLM (Shao et al. 2023) Multi-Turn Artificial Interview
CharacterGLM (Zhou et al. 2023) Multi-Turn Artificial Dialogue
RoleLLM (Wang et al. 2023c) Sentence
CharacterEval (Tu et al. 2024) Sentence
RoleCraft (Tao et al. 2023) Sentence
InCharacter (Wang et al. 2023a) Multi-Turn Artificial Interview
DITTO (Lu et al. 2024) Multi-Turn Automated Dialogue
@@
Table 1: Comparison of Beyond Dialogue with other datasets or framework. “✓” indicates that the dataset/framework meets the given standard or includes the specified dimension in the role profile, while “✗” indicates that it does not. “Sentence” indicates that the dataset/framework evaluates performance based on the final sentence of multi-turn dialogues. Other methods assess performance through dialogues/interviews generated manually or automatically.
@@
General Role-playing Agents. Recent research on general role-playing agents has largely overlooked the issue of bias between predefined profiles and scenario-specific dialogues. Instead, the focus has been on constructing role-playing datasets, primarily through manual curation (Zhou et al. 2023), extraction from novels (Li et al. 2023; Tao et al. 2023), or generation using GPT (Shao et al. 2023; Wang et al. 2023c), to acquire extensive annotated role-playing data. We compared various current Role-Playing Datasets/Frameworks and summarized five key dimensions for constructing role profiles and evaluations in Table 1: Style, Character, Personality, Emotion, and Relationship. Additionally, scenarios provide the background and context for dialogues, making the role’s conversations more contextual and realistic (Chen et al. 2023; Shao et al. 2023).
@@
Role-playing Evaluation. The initial step in evaluating role-playing models involves generating dialogues for assessment, employing three primary methods: Providing Historical Dialogues (Tu et al. 2024; Wang et al. 2023c; Chen et al. 2023; Tao et al. 2023), Manual Dialogue (Zhou et al. 2023; Wang et al. 2023a; Shao et al. 2023), and Automated Dialogue with LLMs (Lu et al. 2024). The first method is limited to evaluating the model within the scope of individual sentences; however, single sentences are likely to fail to fully capture the entire role profile and may be biased by exposure to novels or scripts used during pre-training. In contrast, the second and third methods can generate complete scenario dialogues, which are more conducive to comprehensive evaluations.
@@
There are three common methods for evaluating dialogues generated by models: Metric-based Evaluation (Li et al. 2023; Wang et al. 2023c), Human Evaluation (Zhou et al. 2023), and “LLMs as Judges” (Shao et al. 2023; Wang et al. 2023c; Lu et al. 2024). The first method primarily assesses the model’s retention of the standard response, which is derived from the original text of the novel or manually annotated compilation. The second method, while more accurate, is costly and difficult to replicate, limiting its broader application. The third method, however, is attracting increasing attention due to its high efficiency, low cost, and scalability.
@@
Hybrid Task Training LLMs. In recent years, hybrid task training has emerged as a pivotal method for fine-tuning LLMs. For instance, the LLaMA Writing Assistant is trained on a variety of textual tasks (Wang et al. 2023b). Incorporating tasks like summarization, text generation, and reasoning within reading comprehension has greatly improved the model’s language understanding and QA abilities (Cheng, Huang, and Wei 2023). Furthermore, training with a combination of general and role-specific instructions has improved the model’s role-playing abilities (Tao et al. 2023). These studies demonstrate that incorporating tasks directly or indirectly related to downstream tasks in hybrid training can significantly enhance model performance.
Beyond Dialogue Framework
Figure 2: Illustration of Beyond Dialogue Training Framework and Automated Dialogue Evaluation. The left side shows the training phases, which include role-playing dialogue, chit-chat, and profile alignment. The profile alignment results are utilized to adjust each scenario’s dialogue profiles, eliminating training biases. On the right, the LLM generates random scenarios and roles for multi-turn dialogues with the model, followed by an evaluation using objective questions to obtain quantitative metrics of the model’s role-playing capabilities.
As shown in Figure 2, In the Beyond Dialogue training framework, role profiles are aligned and adjusted for each scenario to create ’pure’ role-playing dialogue data, which is then combined with alignment reasoning and chit-chat data for model training. In the Automated Dialogue Evaluation, the pipeline generates random scenarios and multi-turn dialogues to interact with the model, using objective questions to quantify its role-playing capabilities.
The design and implementation details of each component will be elaborated in the subsequent chapters.
Alignment Dataset Construction
@@
Figure 3: An schematic of our automated pipeline for constructing Beyond Dialogue dataset.
@@
To advance general role-playing, we require a diverse set of role settings and dialogue data. Comparing novel-extracted dialogues (Chen et al. 2023) with those generated by GPT models (Wang et al. 2023c; Li et al. 2023) shows that the latter lacks human likeness, while the former better captures role nuances. However, datasets that meet the requirement for diverse roles, extracted solely from novels, are not open-source due to high costs and significant labor. Therefore, we propose a low-cost (as discussed in §E), fully automated pipeline for constructing a role-playing dialogue dataset (Figure 3) and introduce a method to address biases and fine-grained alignment issues mentioned in the §Introduction:
  1. 1.
@@
    Role-playing Dialogue Dataset Construction: We first segment the text and apply a role frequency threshold to filter chunks. Given the large initial chunk volume, we employ open-source models to extract scenarios and evaluate chunks, retaining only those that reflect role traits and belong to single scenes. Next, we compare various LLMs and ultimately select GPT-4o for dialogue extraction and validation, ensuring the dialogues are coherent and free from conflicts with role profiles. By applying rules and open-source models to filter all chunks, we generate high-quality chunks for proprietary model processing. This approach enables the low-cost, fully automated construction of a high-quality role-playing dialogue dataset, laying a solid foundation for subsequent derivative tasks and model training.
  2. 2.
@@
    Profile Alignment: After constructing the dataset, we use GPT-4o to align each multi-turn dialogue across five dimensions: Character, Style, Emotion, Relationship, and Personality (CSERP). Using an innovative prompting mechanism (as detailed in §F), GPT-4o provides detailed reasoning processes, such as identifying how a dialogue line reflects a specific speaking style. This approach generates fine-grained CSERP alignment tasks, helping the model maintain its reasoning capabilities and enhancing its ability to perceive and adhere to profiles.
  3. 3.
@@
    Profile Adjustment: Based on the alignment results, we dynamically adjust the Profile settings for each dialogue to ensure consistency. This process addresses the common issue of bias between scene dialogues and role profiles in role-playing datasets, aligning the training inputs (Profile prompts) with outputs (dialogue labels), thereby ensuring effective training.
All prompts used for dataset construction and alignment can be found in §F. A detailed explanation of the database construction process is provided in §A, along with the statistical information about our dataset, obtained using this method, which can be found in §A. This section also includes manually collected and organized role profile statistics, covering the five dimensions of CSERP. A discussion of the low cost is provided in §E.
Supervised Finetuning
Using the constructed dataset, we fine-tuned LLMs with a mix of Role-Playing Dialogue data $D_{r}$, Deriving CSERP data $D_{d}$, and Chit-Chat data $D_{c}$ in a 1:5:4 ratio. $D_{r}$ dataset was aligned and adjusted to ensure consistency between role profiles and dialogue across scenarios. The learning objective was to minimize the aggregate negative log-likelihood across all datasets, as described by:
$\min_{\pi}\mathbb{E}_{(s,a)\sim D_{r}\odot D_{d}\odot D_{c}}\left[-\log\pi(a|s)\right]$

### v6

@@ Abstract
The rapid advancement of large language models (LLMs) has revolutionized role-playing, enabling the development of general role-playing models. However, current role-playing training has two significant issues: (I) Using a predefined role profile to prompt dialogue training for specific scenarios usually leads to biases and even conflicts between the dialogue and the profile, resulting in training biases. (II) Models learn to imitate the role based solely on the profile, neglecting profile-dialogue alignment at the sentence level. To overcome the aforementioned hurdles, we propose a novel framework Beyond Dialogue, which introduces “beyond dialogue” tasks to align dialogue with profile traits for each scenario, eliminating biases during training. Furthermore, the framework achieves a sentence-level fine-grained alignment between profile and dialogue through an innovative prompting mechanism that generates reasoning data for training. Moreover, the aforementioned methods are fully automated and low-cost. Experimental results demonstrate our model excels in adhering to role profiles, outperforming most proprietary general and specialized role-playing baselines. The code and data are provided in https://github.com/yuyouyu32/BeyondDialogue.
Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model
Yeyong Yu1, Runsheng Yu3, Haojie Wei2, Zhanqiu Zhang2, Quan Qian*1 1School of Computer Engineering & Science, Shanghai University 2LIGHTSPEED 3The Hong Kong University of Science and Technology Correspondence: yuyeyong@shu.edu.cnqqian@shu.edu.cn
1 Introduction
The rapid advancement of large language models (LLMs) has demonstrated their significant potential in user interactions (Achiam et al., 2023; Yang et al., 2024; Dubey et al., 2024). A particularly promising area is the development of role-playing LLM agents, which are capable of simulating both real and fictional roles to deliver immersive and interactive experiences (Chen et al., 2024b, a). Applications like Character AI [^1] https://character.ai, Replika  [^2] https://replika.com, and Baichuan-NPC (Yang et al., 2023) exemplify the growing impact of these agents. However, the proprietary nature of these platforms restricts users’ ability to customize specific role models. Thus, the development of an open-source, general role-playing model is imperative.
@@ 1 Introduction
At the current stage of role-playing dialogue training tasks, role profiles are typically manually compiled (Shao et al., 2023) or generated through LLM summaries (Zhou et al., 2024), including information such as personality and speaking style as depicted throughout a whole novel or script (Zhou et al., 2024; Li et al., 2023). However, a bias exists between the training dialogues extracted from individual scenarios and the predefined role profiles. As an illustration in Figure 1, Hermione’s speaking style is predefined into four categories, which are manually extracted from the entire content of the novel. However, the training dialogues in Figure 1 predominantly reflect only two styles (“bookish–blue” and “encouraging–green”). The bias arises because a single scenario often fails to fully capture the predefined profile, resulting in misalignment even contradiction between the prompted profile and the training dialogues. Such biases are common: from the HPD dataset published by Chen et al. (2023), we sampled 800 scenarios and found that 83.2% of the dialogues exhibited bias in relation to the predefined role profiles, where at least one dimension, such as character or speaking style, was unaligned. As shown in our § 4.2 experiments, biases between predefined role profiles and the training dialogue corpus can mislead LLMs, impairing its ability to accurately reflect the role profile in dialogues. Meanwhile, the model has only learned a vague mapping from profile to dialogue, without fine-grained alignment of specific traits to their corresponding dialogues’ sentence (as indicated by the red arrow in Figure 1). Consequently, the model is unable to learn how “personality traits are revealed in dialogues”.
@@
We propose a new training role-playing framework — Beyond Dialogue to overcome the aforementioned issues and advance toward general role-playing. We incorporate powerful LLM through a prompting mechanism approach to align role profiles with scenario dialogues. This not only ensures the reliability of the LLM’s reasoning but also explicitly creates a fine-grained alignment dataset at the sentence level. Taking inspiration from actors learning to play different roles — understanding the performance of various role traits in scenarios to enhance their portrayal (Zarrilli, 2005) — we also employ fine-grained alignment tasks to train the role-playing model. Compared to previous methods that relied on profiles to loosely guide dialogue generation (Lu et al., 2024; Tang et al., 2024), these alignment tasks directly link profile attributes to dialogue sentences, allowing the model to better capture role traits at the sentence level and improve its effectiveness in role portrayal.
Traditional evaluation methods rely on subjective assessments (e.g., human ratings or LLM-based scoring), leading to inconsistent outcomes (Wang et al., 2024a; Tu et al., 2024). Our approach converts evaluation tasks into objective ones (e.g., multiple-choice and judgment questions), clearly defining the model’s ability to follow role profiles as the evaluation criterion. Combined with “LLMs as Judges” (Kim et al., 2023), this approach objectively quantifies role-following capabilities with measurable criteria as demonstrated in § B.1.
We applied Beyond Dialogue to bilingual chat LLM baselines — Qwen2 (Yang et al., 2024) and Mistral-Nemo (MistralAI, 2024) — to evaluate its effectiveness. Extensive experiments demonstrate that Beyond Dialogue significantly improves LLMs’ ability to follow and portray role profiles. Under our framework, the trained models outperformed advanced general baselines like GPT-4o (Hurst et al., 2024) and specialized role-playing baselines like Baichuan-NPC-Turbo (Yang et al., 2023) in key aspects of role-playing performance.
@@
    We identified a bias in role-playing training where the predefined role profile misaligns with the dialogues in a specific scenario. Furthermore, we demonstrated that this bias significantly hinders the model’s ability to effectively follow the predefined profile.
@@
    We propose Beyond Dialogue, a novel general role-playing training framework that specifically addresses the identified bias by using LLMs with a prompting mechanism to align role profiles with scenario-specific dialogues. This approach generates fine-grained alignment tasks at the sentence level, which are integrated into the training, further enhancing the effectiveness of role portrayal.
@@
    We introduce a novel evaluation pipeline that converts all assessment tasks into objective tasks, focusing on the model’s ability to follow user-defined role profiles. By combining automatic dialogues with the “LLMs as Judges” method, our approach ensures both efficiency and reproducibility in role-playing LLMs evaluation.
@@
LIGHT (Urbanek et al., 2019) Sentence
PDP (Han et al., 2022) Sentence
ChatHaruhi (Li et al., 2023) Sentence
HPD (Chen et al., 2023) Sentence
CharacterLLM (Shao et al., 2023) Multi-Turn Artificial Interview
CharacterGLM (Zhou et al., 2024) Multi-Turn Artificial Dialogue
RoleLLM (Wang et al., 2024a) Sentence
CharacterEval (Tu et al., 2024) Sentence
RoleCraft (Tao et al., 2024) Sentence
InCharacter (Wang et al., 2024b) Multi-Turn Artificial Interview
DITTO (Lu et al., 2024) Multi-Turn Automated Dialogue
@@
Table 1: Comparison of Beyond Dialogue with other datasets or framework. “✓” indicates that the dataset/framework meets the given standard or includes the specified dimension in the role profile, while “✗” indicates that it does not. “Sentence” refers to evaluation based on the final sentence of multi-turn dialogues, while other methods assess performance through manually or automatically generated dialogues/interviews.
2 Related Work
@@
General Role-playing Agents. Recent research on general role-playing agents has largely overlooked the issue of bias between predefined profiles and scenario-specific dialogues. Instead, the focus has been on constructing role-playing datasets, primarily through manual curation (Zhou et al., 2024), extraction from novels (Li et al., 2023; Tao et al., 2024), or generation using GPT (Shao et al., 2023; Wang et al., 2024a, 2025), to acquire extensive annotated role-playing data. We compared various current Role-Playing Datasets/Frameworks and summarized five key dimensions for constructing role profiles and evaluations in Table 1: Style, Character, Personality, Emotion, and Relationship. Additionally, scenarios provide the background and context for dialogues, making the role’s conversations more contextual and realistic (Chen et al., 2023; Shao et al., 2023).
@@
Role-playing Evaluation. The initial step in evaluating role-playing models involves generating dialogues for assessment, employing three primary methods: Providing Historical Dialogues (Tu et al., 2024; Wang et al., 2024a; Chen et al., 2023; Tao et al., 2024), Manual Dialogue (Zhou et al., 2024; Wang et al., 2024b; Shao et al., 2023), and Automated Dialogue with LLMs (Lu et al., 2024). The first method is limited to evaluating the model within the scope of individual sentences; however, single sentences are likely to fail to fully capture the entire role profile and may be biased by exposure to novels or scripts used during pre-training. In contrast, the second and third methods can generate complete scenario dialogues, which are more conducive to comprehensive evaluations.
@@
There are three common methods for evaluating dialogues generated by models: Metric-based Evaluation (Li et al., 2023; Wang et al., 2024a), Human Evaluation (Zhou et al., 2024), and “LLMs as Judges” (Shao et al., 2023; Wang et al., 2024a; Lu et al., 2024). The first method primarily assesses the model’s retention of the standard response, which is derived from the original text of the novel or manually annotated compilation. The second method, while more accurate, is costly and difficult to replicate, limiting its broader application. The third method, however, is attracting increasing attention due to its high efficiency, low cost, and scalability.
@@
Hybrid Task Training LLMs. In recent years, hybrid task training has emerged as a pivotal method for fine-tuning LLMs. For instance, the LLaMA Writing Assistant is trained on a variety of textual tasks (Wang et al., 2023). Incorporating tasks like summarization, text generation, and reasoning within reading comprehension has greatly improved the model’s language understanding and QA abilities (Cheng et al., 2023). Furthermore, training with a combination of general and role-specific instructions has improved the model’s role-playing abilities (Tao et al., 2024). These studies demonstrate that incorporating tasks directly or indirectly related to downstream tasks in hybrid training can significantly enhance model performance.
3 Beyond Dialogue Framework
@@
Figure 2: A schematic of our automated pipeline for constructing coarse Role-playing dialogue dataset.
Figure 3: Illustration of Refined Dataset Construction, Supervised Finetuning, and Automated Dialogue Evaluation Framework. The left side shows the Refined Dataset Construction and Supervised Finetuning phases, the profile alignment results are utilized to adjust each scenario’s dialogue profiles, eliminating training biases. On the right, the LLM generates random scenarios and roles for multi-turn dialogues with the model, followed by an evaluation using objective questions to obtain quantitative metrics of the model’s role-playing capabilities.
We propose the Beyond Dialogue Framework for role-playing, which includes three key parts: i) in the Alignment Dataset Construction stage, role profiles are aligned and adjusted for each scenario to create “pure” role-playing dialogue data. ii) in the Supervised Finetuning stage, the “pure” role-playing dialogue data is combined with alignment reasoning and chit-chat data for model training. iii) in the Automated Dialogue Evaluation, the pipeline generates random scenarios and multi-turn dialogues to interact with the model, using objective questions to quantify its role-playing capabilities.
Details of each component’s implementation will be provided in the following chapters.
3.1 Alignment Dataset Construction
@@
To advance general role-playing, we require a diverse set of role settings and dialogue data. However, the dialogues generated by GPT models (Wang et al., 2024a; Li et al., 2023) lacks human likeness. The novel-extracted dialogues (Chen et al., 2023) better captures role nuances but these datasets are expensive and labor-intensive to produce, making them less accessible. Therefore, we propose a low-cost (as shown in § E), fully automated pipeline (Figure 2) for constructing a role-playing dialogue dataset and introduce a method to address biases and fine-grained alignment issues.
3.1.1 Coarse Role-playing Dialogue Dataset Construction
@@
We first segment the text and apply a role frequency threshold to filter chunks (Figure 2.1). Given the large initial chunk volume, we employ open-source models to extract dialogue scenarios and evaluate chunks, retaining only those that reflect role traits and belong to single scenes (Figure 2.2). Next, through systematic comparison of multiple LLMs, we select GPT-4o for its optimal performance in dialogue extraction and scene reconstruction (see Figure 2 panels 3-4, with accuracy metrics and cost analysis in Table 5 of § A.2), while implementing rigorous verification to ensure dialogue coherence and profile consistency. This pipeline (with implementation details in § A.1) enables the low-cost, fully automated construction of a high-quality role-playing dialogue dataset.
3.1.2 Refined Dataset Construction by Profile Alignment and Profile Adjustment
@@
Profile Alignment: After dialogue dataset construction, we use GPT-4o to align each multi-turn dialogue across five dimensions: Character, Style, Emotion, Relationship, and Personality (CSERP) as shown in Figure 2.5. Through our innovative prompting mechanism (outlined in § F.2), LLMs are able to analyze dialogues through step-by-step reasoning, explicitly linking each sentence to corresponding profile traits, thereby achieving sentence-level alignment (see § G.2 for examples). This approach generates fine-grained CSERP alignment tasks, helping the model maintain its reasoning capabilities and enhancing its ability to perceive and adhere to profiles during training. All prompts used for dataset construction and alignment are provided in § F.1 and § F.2, respectively.
@@
Profile Adjustment: Based on the alignment results, we dynamically adjust the profile settings for each dialogue to ensure consistency, as indicated by the red arrow in Figure 3. Specifically, we refine the predefined profile by removing traits that are not reflected in the dialogue while incorporating scenario-relevant Emotional and Relational attributes. This addresses the common issue of bias between scene dialogues and role profiles in role-playing datasets, aligning the training inputs (profile prompts) with outputs (dialogue labels), thereby ensuring effective training.
3.2 Supervised Finetuning
Using the constructed dataset, we fine-tuned LLMs with a mix of Aligned Role-Playing Dialogue data $D_{r}$, deriving Alignment CSERP data $D_{a}$, and Chit-Chat data $D_{c}$ (sourced from open-source chit-chat datasets) in a 1:5:4 ratio. Initially, we adopted a 1 ($D_{r}$) : 9 ($D_{c}$) ratio, following Tao et al. (2024), and subsequently replaced part of $D_{c}$ with alignment data $D_{a}$ (since each $D_{r}$ generates five $D_{a}$ samples) to better capture role-specific features. The $D_{r}$ dataset ensures that model’s output dialogue are aligned with corresponding role profile, and $D_{a}$ focuses on explicit alignment tasks to refine profile-dialogue relationships. Meanwhile, $D_{c}$ preserves the model’s general conversational ability. The learning objective was to minimize the aggregate negative log-likelihood across all datasets, as described by:
$\min\nolimits_{\pi}\mathbb{E}_{(s,a)\sim D_{r}\odot D_{a}\odot D_{c}}\left[-\log\pi(a|s)\right]$